PLEASE HELP I AM STUCK WITH THIS PROBLEM
(OF THE FOLLOWING,WHICH IS THE MOST NEARLY EQUAL TO 2/3?
(A)3/4
(B)5/6
(C)7/9
(D)11/15
(E)15/21
PLEASE DESCRIBE ME TOO HOW YOU DID?

Answer by rfer(16322)

2/3=0.667
------------
3/4=0.75
5/6=0.83
7/9=0.778
11/15=0.733
15/21=0.71
you decide